Ian goes to the grocery store after work. He buys 3/4 pounds of oranges at $0.69 a pound, 1 1/2 pounds of apples at $0.89 a pound, and 3 1/4 pounds of bananas at $0.39 a pound. He also buys 2 gallons of milk at $2.19 a gallon. Ian gives the cashier a ten-dollar bill. How much change does ian recieve
Answer by drk(1908)

oranges: (3/4)*(0.69) = . . . . . .(0.5175)
apples: 1.5*(0.89) = . . . . . . . .(1.335)
bananas: 3.25*(0.39) = . . . . . .(1.2675)
milk: 2*(2.19) = . . . . . . . . . . .(4.38)
This is a total amount of $7.50
since he gave the cashier a ten dollar bill, he will get $2.50 back in change